OpenStudy (anonymous):

I hope someone can help me. I have to solve the following rational expression: 5 - a-1/a+2 = a^2 - 1/a + 2 (the slashes represent fractions). I am coming up with an answer of 4a + 9/ a+2 = a - 1 but it doesn't seem right. Please help.

OpenStudy (amistre64):

cross multiply then combine lke terms

OpenStudy (amistre64):

\[\frac{5-a-1}{a+2} = \frac{a^2-1}{a+2}\] that -1 looks suspicious

OpenStudy (amistre64):

since the denoms are the same, just ignore them... equate the tops

OpenStudy (anonymous):

Make the 5 a fraction of 5/1 correct? If so, do I need to change that to the denominator a + 2 or just include the 5 into the answer?